OpenStudy (anonymous):

write each precent as a fraction in simplest form and a decimal 225% and 201% write each precent as a fraction in simplest form and a decimal 225% and 201% @Mathematics

OpenStudy (anonymous):

225% = 2.25 because percents are just numbers multiplied by 100, so 225%/100% = 2.25 Using 2.25 to get a fraction, we know that 2.25 = 2 whole + (1/4) = 9/4

OpenStudy (anonymous):

Following this logic, 201% = 2.01 = 201/100